2. As you are leaving the community, you notice another new community just down the street. Because you are in the area, you decide to inquire about it.
a) The sales representative here tells you they also have two floor plans available, but they only have 38 homes available. Write an equation that illustrates the situation. Use x and y to denote floor plan #1 and floor plan #2 respectively.
b) The representative tells you that floor plan #1 sells for $175,000 and floor plan #2 sells for $200,000. She also mentions that all the available houses combined are worth $7,200,000. Write an equation that illustrates this situation. Use the same variables you used in part a.
c) Use elimination to determine how many houses with each floor plan are available. Explain how you arrived at your answer.

I am totally clueless. I am familiar with word problems, and the elimination method. I have a major problem attempting to figure out how to translate the given information into an equation. I cannot solve for something that I cannot first translate. PLEASE HELP!!!

Answer by stanbon(75887)

As you are leaving the community, you notice another new community just down the street. Because you are in the area, you decide to inquire about it.
-----------------------------
a) The sales representative here tells you they also have two floor plans available, but they only have 38 homes available. Write an equation that illustrates the situation. Use x and y to denote floor plan #1 and floor plan #2 respectively.
Quantity Equation: x + y = 38
-----------------------------
b) The representative tells you that floor plan #1 sells for $175,000 and floor plan #2 sells for $200,000. She also mentions that all the available houses combined are worth $7,200,000. Write an equation that illustrates this situation. Use the same variables you used in part a.
Value Equation: 175000x + 200,000y = 7200000
-------------------------------------------------------
c) Use elimination to determine how many houses with each floor plan are available. Explain how you arrived at your answer.
---
Multiply thru the Quantity equation by 175000 to get:
175000x + 175000y = 38*175000
---
Subtract that equation from the value equation and solve for "y":
25000y = 720000-38*175000
y = 22 (# of #2 floor-plan houses)
x = 38-y = 38-22 = 16 (# of #1 floor-plan houses)
======================================================
Comment: You many not realise it but these word problems are
attempting to get you to a new level of reading ability. It
would be like me trying to read a medical text---I would find
out that "I cannot read".
======================================================
Cheers,
Stan H.